Two high speed trains are 270 miles apart and traveling toward each other. they meet in 2 hours. if one train's speed is 10 miles per hour faster than the other, find the speed of each train
The speed of the slower train is? mph. (Round to two decimal places as needed.)

the speed of the faster train is? mph. (Round to two decimal places as needed.)

Answer by TimothyLamb(4379)

---
s = d / t
s = 270 / 2
s = 135 mph
---
x = speed of trainA
y = speed of trainB
s = x + y
y = x + 10
---
s = x + x + 10
s = 2x + 10
---
2x + 10 = 135
2x = 125
---
x = 62.5
y = 72.5
---
answer:
speed of trainA = 62.5 mph
speed of trainB = 72.5 mph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php